# Picogram-Level Nanoplastic Analysis with Nanoelectromechanical System Fourier Transform Infrared Spectroscopy: NEMS-FTIR

### v2

@@ Abstract
We present a photothermal infrared spectroscopy-based approach for the chemical characterization and quantification of nanoplastics. By combining the high sensitivity of nanoelectromechanical systems (NEMS) with the wide spectral range and ubiquity of commercially available Fourier transform infrared (FTIR) spectrometers, NEMS–FTIR offers a cost-efficient and cryogen-free option for the rapid, routine analysis of nanoplastics in aqueous samples. Polypropylene, polystyrene, and polyvinyl chloride nanoplastics with nominal diameters ranging from 54 to 262 nm were analyzed by NEMS–FTIR with limits of detection ranging from 102 pg to 355 pg. The absorptance measured by NEMS–FTIR could be further converted to absolute sample mass using the attenuation coefficient, as demonstrated for polystyrene. Thanks to the wide spectral range of NEMS-–FTIR, nanoplastic particles from different polymers could be readily identified, even when present in a mixture. The potential of NEMS–FTIR for the analysis of real samples was demonstrated by identifying the presence of nanoplastics released in water during tea brewing. Polyamide leachates were identified in the brewing water without sample pre–concentration. Accelerated aging of the tea bags under elevated temperature and UV radiation showed further release of polyamide over time.
Keywords: nanoplastics, nanoparticles, NEMS–FTIR, Fourier transform infrared spectroscopy, photothermal sensing, nanoelectromechanical system, nanomechanical resonator, tea bag.
@@ INTRODUCTION
Nanoplastics have become ubiquitous and pose significant environmental and health risks due to their high reactivity, potential to transport pollutants, and widespread dispersion across ecosystems [1, 2, 3, 4, 5, 6, 7]. The ability of nanoplastics to penetrate deep into tissues underscores the need for their routine chemical characterization and monitoring [8]. Sampling of nanoplastic particles is particularly challenging due to their nanoscale dimensions and low concentrations, which often exceed the capabilities of many analytical techniques [9, 10, 11]. While state–of–the–art techniques allow for routine microplastic detection and identification, nanoplastics, on the other hand, remain a challenge to this day.
@@
Mass spectrometry–based techniques can be used for the bulk chemical identification of nanoplastics [10, 11, 12]. Pyrolysis–gas chromatography–mass spectrometry (Py–GC/MS) has been extensively applied to detect various types of nanoplastics, including polystyrene (PS), polypropylene (PP), polycarbonate (PC), polyethylene terephthalate (PET), polymethyl methacrylate (PMMA), polyvinyl chloride (PVC), polyethylene (PE), and Nylon 6/66, with reported limits of detection (LoDs) as low as 1 ng/L [13, 14]. Using the same technique, Li et al. [15] estimated an LoD of 10 ng for two polymer types of nanoplastics in water, PS, and PMMA. A combination of atomic force microscope–IR (AFM–IR) spectroscopy and Py–GC/MS was also used to detect and quantify PE and PVC particles (20-–1000 nm), with LoDs of 0.28 g for PE and 1.54 g for PVC [16]. Alternative approaches, such as thermal desorption–proton transfer reaction-mass spectrometry (TD–PTR–MS), demonstrated remarkable sensitivity, with Materić et al. [17] detecting pure 1 m PS particles with an LoD of 0.34 ng, allowing detection of these particles in environmental samples as small as 1 mL without preconcentration. While mass spectrometry–based techniques provide excellent capabilities, their complexity and cost hinder their use in routine environmental monitoring [11, 12, 18].
@@
FTIR spectroscopy with its broad spectral range allows functional group chemical identification of unknown samples and is one of the most extensively used methods in microplastic analysis [19, 20]. Commonly used attenuated total reflectance–FTIR (ATR–FTIR) spectroscopy is a contact bulk analysis technique with a detection limit in the order of 10 ng [21]. FTIR microscopy methods, such as µ–FTIR scanning microscopy and focal plane array–FTIR (FPA–FTIR) microscopy, offer a contact–less FTIR solution, reaching LoDs as low as 450 pg when used with a cryogenically cooled MCT array [22]. However, diffraction limits the reliable analysis of single particles with sizes below 10 m [19, 20].
@@
Quantum cascade laser infrared (QCL–IR) microscopy provides higher sensitivity, albeit with a narrower spectral range and additional issues with coherence artifacts. Although QCLs cover the fingerprint region relevant for polymer identification and provide stronger spectral signals due to their high source intensity, they typically have higher noise levels and limited spectral coverage, missing informative regions such as the C–-H stretching vibrations around 3000 cm-1 and the low-wavenumber region around 700 cm-1. Full–range spectra, as enabled by FTIR, also allow for chemometric analysis and spectral deconvolution, which can help identify and quantify nanoplastics in multicomponent samples. While such methods can also work within narrower spectral windows, having access to broader spectral information can make differentiation easier and more reliable. QCL–IR microscopy is a complementary method to FTIR microscopy and is therefore sometimes offered in tandem.222HYPERION II from Bruker Corporation All standard IR spectroscopy techniques suffer from Mie scattering–related spectral artifacts when measuring particulate samples.[23]
@@
Scattering, diffraction, and coherence issues can be mitigated by QCL–based photothermal scanning methods, such as optical–-photothermal IR (O-PTIR) spectroscopy or AFM-IR microscopy. O–PTIR has proven effective for the imaging with sub–micrometer resolution of individual PS nanoparticles as small as 250 nm in mammalian tissues [24] and identification of plastics (> 600 nm) containing polysiloxanes and imides, originating from the steam disinfection of silicone baby teats [25]. Even better lateral resolution and single nanoparticle analysis can be obtained by AFM-IR [26]. However, AFM–IR faces challenges with low signal–to–noise ratio (SNR) and slow imaging speed [19].
@@
Raman spectroscopy has the advantage of probing at visible wavelengths with low interference from water and, similar to O–PTIR, allows for the analysis of plastic particles down to sub-micrometer sizes, with spatial resolution sufficient to resolve and chemically identify individual particles. [27]. Surface–enhanced Raman spectroscopy (SERS) has emerged as a promising technique. While Zhou et al. [28] and Hu et al. [29] reported LoDs as low as 5 g/mL for PS particles as small as 50 nm using conventional SERS substrates, integration with sample preparation methods has pushed the sensitivity down to g/mL [30]. Hyperspectral stimulated Raman scattering (SRS) microscopy has been demonstrated as a powerful tool for the chemical imaging of nanoplastics, enabling identification and classification of particles down to 130 nm [31]. However, SERS suffers from issues such as spectral artifacts and reliance on carefully engineered substrates, limiting its versatility [11, 32]. SRS microscopy, despite offering rapid detection and high sensitivity, relies on complex and expensive instrumentation [33].
@@
Although existing analytical techniques such as Py–GC/MS, QCL–IR, O–PTIR, or SERS offer high sensitivity and specificity, they remain impractical for routine environmental monitoring, particularly for water suppliers and official control laboratories [34].
@@
Nanoelectromechanical systems IR spectroscopy (NEMS–IR) is a technique based on the photothermal effect that has been introduced in recent years. It employs NEMS membrane resonators, or chips, as its key component [35, 36]. The NEMS chip can be used simultaneously as a sample carrier and detector. The NEMS chips, typically consisting of a pre–stressed few–nanometer–thin silicon nitride (SiN) membrane, are robust and allow for a wide variety of sampling techniques [37], including direct aerosol collection [38, 36, 35] and drop casting. When a sample deposited on the NEMS chip’s surface absorbs IR light, local heating occurs, causing thermal expansion and tensile stress reduction. The corresponding frequency detuning of the NEMS chip is proportional to the absorbed power, and the frequency shift can be monitored using a closed–loop oscillation scheme in combination with a frequency counter, as demonstrated by Bešić et al. [39, 40]. NEMS–IR has shown picogram sensitivity at room temperature over a large spectral range, from the ultraviolet to the far–IR [37]. This technique has been used to analyze polymer nanoparticles [41, 42], pharmaceutical compounds [36, 43], polymer micelles [44], thin polymer films [21] and explosives [45]. NEMS–IR with in–situ thermal desorption (TD) analysis has also been demonstrated for separating simple analyte mixtures by Luhmann et al. [35].
@@
NEMS–IR has only been used for dispersive spectroscopy, typically relying on narrow-band tunable QCLs. This paper introduces NEMS–FTIR spectroscopy (see Figure 1) as a new approach to the chemical characterization and quantification of nanoparticles, demonstrated here with nanoplastics. By combining the broad spectral coverage, wide availability, and ease of use of commercially available FTIR spectrometers with the high sensitivity of nanomechanical photothermal detection using NEMS membrane resonators, NEMS-–FTIR enables artifact–free measurements with picogram–level detection limits — all without the need for cryogenic cooling. Thanks to its straightforward sampling of nanoparticles dispersions directly onto the disposable NEMS chip, NEMS-–FTIR provides a powerful bulk analysis method for the routine detection and characterization of nanoplastics. Optomechanical metalized cantilevers have previously been used as photothermal detectors in combination with FTIR sources [46], but their lower sensitivity [47], reliance on optical readout, and limited surface area for sample collection have constrained their practical application.
@@
Here we demonstrate the capability of NEMS–FTIR for the sampling, identification, and quantification of nanoplastics dispersions and nylon-based polyamide (PA) tea bag leachates. Many studies have analyzed the release of micro- and nanoplastic particles from tea bags during tea preparation using different analytical methods such as scanning transmission X–ray microscopy [48], IR photothermal heterodyne microscopy [49], Raman imaging [50], and ATR–FTIR spectroscopy [51, 52]. Most of these, however, require a much higher tea bag–to–water ratio than that applied in this study [52, 51] or involve sample up–concentration steps such as ultracentrifugation [51], density separation [48], or evaporation drying to obtain a powder [52]. In contrast, we demonstrate here that NEMS-–FTIR can chemically identify PA leachates in the brewing water from a single tea bag. NEMS-FTIR analysis of the leachate required just a few nL sample aliquot and did not require any pre–concentration steps.
@@
Aqueous dispersions of three different polymers, PS (nominal diameter: 100 nm), PP (nominal diameter: 54 nm), and PVC (nominal diameter: 262 nm), were deposited via drop casting on individual NEMS chips. Figure 2A–C shows PS, PP, and PVC nanoparticles after deposition onto the central circular perforation of the NEMS chips (see Figure 1A). These scanning electron microscopy (SEM) images highlight the generally spherical morphology and relatively uniform size distributions of the nanoplastic particles studied. However, there are areas of nanoparticle agglomerations on the membrane of the NEMS chip caused by the coffee ring effect during droplet drying (see Figure S1 in Supporting Information). After drying, nanoparticles were confined within the perforated area of all analyzed NEMS chips, except in a few instances where minor droplet spreading occurred beyond this region. The spread–out particles still remained located within the confines of the NEMS membrane, which is the active area relevant for signal detection.
Spectral data were obtained by depositing a 20 nL drop of various nanoplastic concentrations on the NEMS chips (N=3 for each concentration). The NEMS–FTIR spectra of PS nanoplastic particles for each concentration, power–corrected, normalized using the SiN peak at 835 cm-1, and converted into absorbance, are shown in Figure 2D. The broad SiN peak, originating from the membrane material and spanning the 1000–700 cm-1 region, serves as a useful internal standard for spectral normalization. The details of the spectral processing procedure are described in "Processing of NEMS–FTIR spectra" in the Materials and Methods section. The spectra of PP and PVC nanoparticles are available in the Supporting Information in Figures S2 and S3, respectively.
@@
The IR peaks in the NEMS–FTIR spectra of PS nanoparticles are in agreement with data reported in the literature [53, 54, 55, 56]. Notably, PS is characterized by peaks at 753 and 700 , as well as by a mono–substituted benzene ring, as evidenced by a series of weak overtone and combination bands between 2000 and 1650 (so–called ’benzene fingers’), often difficult to detect in the spectrum if the concentration of molecules containing the benzene ring is low [55]. However, these peaks are clearly visible in the NEMS–FTIR spectra of PS, even for the very low masses in the nanogram range.
The NEMS–FTIR spectra of PP and PVC nanoparticles also display peaks which are in agreement with prior studies using FTIR spectroscopy [55, 57], as well as with the measurements obtained using early IR spectrophotometers employing prisms [58]. NEMS–FTIR does not suffer from the red shift typically associated with measurements performed by ATR–FTIR, as observed in previous studies [59]. In ATR–FTIR, slight shifts in some peak positions are observed due to the interaction of the evanescent wave with the sample surface, which depends on both the refractive index and wavelength. The wavelength-dependent penetration depth inherent to ATR measurements can also cause lower–wavenumber bands to appear more intense compared to transmission (or absorbance) spectra. [60, 61, 62, 63, 64]. This artefact is also absent from NEMS–FTIR spectra. ATR–FTIR performance is inherently constrained by the choice of crystal material, as different sample types require different ATR crystals for optimal spectral response [63]. This issue also does not apply to NEMS-FTIR, which relies on direct absorption. NEMS–FTIR generates spectra comparable to those obtained with transmission–FTIR spectroscopy, allowing seamless integration with existing spectral libraries and chemometric tools for efficient and accurate analysis. Furthermore, since the nanoplastic particles are analyzed under high vacuum, the resulting spectra are free from spectral interferences from water molecules and carbon dioxide (CO2).
Figure 2E shows the calibration curves for PS, PP, and PVC nanoparticles. Calibration curves were constructed using the characteristic peaks at 1452 for PS (see inset in Figure 2D), 1377 for PP (see inset in Figure S2), and 1427 for PVC (see inset in Figure S3) and the results showed good linearity (). LoDs were calculated from the calibration curves using the relation , where is the standard deviation of method blanks () and is the slope of the corresponding calibration curve. The results are summarized in Table 1. PP exhibited the steepest calibration slope, resulting from the strongest IR band used for the calibration curve, and consequently had the lowest LoD. Detection limits are lower than those typically achieved by Py–GC/MS [15] and are comparable to the performance of state–of–the–art TD–PTR–MS techniques [17]. The variations in detection limits for different nanoplastic particles did not correlate with nanoparticle size. In the measurements conducted on the chips with higher loadings, the observed nanoparticle agglomerations did not result in deviations in linearity.
@@
The mass of PS nanoparticles present on the NEMS chips was estimated using the attenuation coefficient of PS and the absorptances calculated from the NEMS–FTIR spectra at . The mass estimation procedure is detailed in the "Mass estimation" in the Methods section. The calculation of the sample absorptance accounted for the spatial distribution of the particles, which predominantly formed a circular patch in the center of the perforation on the NEMS chips. Figure 3A shows the correlation between the deposited (amount of deposited PS particles) and estimated (from the measured absorptances of PS particles) mass of nanoplastic particles. The uncertainty on the sample mass deposited on the NEMS chip was estimated from the uncertainty on the tools used for sample preparation and for sample deposition. The largest uncertainty on the sample mass as estimated from the NEMS–FTIR spectra arises from the variability in the measured absorptance . Exact values for the masses deposited on the NEMS chips and estimated masses as calculated from the NEMS–FTIR spectra, along with their standard deviations, can be found in Table S1 in the Supporting Information.
Figure 3B presents a comparison between the NEMS-FTIR spectrum of 20 ng of PS, expressed in absorbance, and a reference spectrum calculated from refractive index data [65] for an equal sample mass and spot size, showing agreement for both peak positions and intensity.

### v3

@@ Abstract
We present a photothermal infrared spectroscopy-based approach for the chemical characterization and quantification of nanoplastics. By combining the high sensitivity of nanoelectromechanical systems (NEMS) with the wide spectral range and ubiquity of commercially available Fourier transform infrared (FTIR) spectrometers, NEMS-FTIR offers a time-efficient and cryogen-free option for the rapid, routine analysis of nanoplastics in aqueous samples. Polypropylene, polystyrene, and polyvinyl chloride nanoplastics with nominal diameters ranging from 54 to 262 nm were analyzed by NEMS-FTIR with limits of detection ranging from 101 pg to 353 pg, one order of magnitude lower than values reported for pyrolysis–gas chromatography–mass spectrometry of nanoplastics. The absorptance measured by NEMS-FTIR could be further converted to absolute sample mass using the attenuation coefficient, as demonstrated for polystyrene. Thanks to the wide spectral range of NEMS–FTIR, nanoplastic particles from different polymers could be readily identified, even when present in a mixture. The potential of NEMS-FTIR for the analysis of real samples was demonstrated by identifying the presence of nanoplastics released in water during tea brewing. Polyamide leachates in the form of fragments and smaller oligomers could be identified in the brewing water without sample pre-concentration, even in the presence of an organic matrix. Accelerated aging of the nylon teabags under elevated temperature and UV radiation showed further release of polyamide over time.
Keywords: nanoplastics, nanoparticles, NEMS-FTIR, Fourier transform infrared spectroscopy, photothermal sensing, nanoelectromechanical system, nylon teabag.
@@ INTRODUCTION
Nanoplastics have become ubiquitous and pose significant environmental and health risks due to their high reactivity, potential to transport pollutants, and widespread dispersion across ecosystems. [1, 2, 3, 4, 5, 6, 7] The ability of nanoplastics to penetrate deep into tissues underscores the need for their routine chemical characterization and monitoring. [8] Sampling of nanoplastic particles is particularly challenging due to their nanoscale dimensions and low concentrations, which often exceed the capabilities of many analytical techniques. [9, 10, 11] While state-of-the-art techniques allow for routine microplastic detection and identification, nanoplastics, on the other hand, remain a challenge to this day.
@@
Mass spectrometry-based techniques can be used for the bulk chemical identification of nanoplastics. [10, 11, 12] Pyrolysis–gas chromatography–mass spectrometry (Py-GC/MS) has been extensively applied to detect various types of nanoplastics, including polystyrene (PS), polypropylene (PP), polycarbonate (PC), polyethylene terephthalate (PET), polymethyl methacrylate (PMMA), polyvinyl chloride (PVC), polyethylene (PE), and Nylon 6/66, with reported limits of detection (LoDs) as low as 1 ng/L. [13, 14, 15] Several studies report Py-GC/MS LoDs in the low-nanogram range for polymer analysis in aqueous samples. [16, 17, 18] A combination of atomic force microscope-IR (AFM-IR) spectroscopy and Py-GC/MS was also used to detect and quantify PE and PVC particles (20–1000 nm), with LoDs of 280 ng for PE and 1.54 g for PVC. [19] Alternative approaches, such as thermal desorption-proton transfer reaction-mass spectrometry (TD-PTR-MS), demonstrated remarkable sensitivity, with Materić et al. [20] detecting pure 1 m PS particles with an LoD of 340 pg, allowing detection of these particles in environmental samples as small as 1 mL without preconcentration. While mass spectrometry-based techniques provide excellent capabilities, their complexity and cost hinder their use in routine environmental monitoring. [11, 12, 21, 15, 22]
@@
FTIR spectroscopy with its broad spectral range allows functional group chemical identification of unknown samples and is one of the most extensively used methods in microplastic analysis. [23, 24] Commonly used attenuated total reflectance - FTIR (ATR-FTIR) spectroscopy is a contact bulk analysis technique with a detection limit in the order of 10 ng. [25] FTIR microscopy methods, such as µ-FTIR scanning microscopy and focal plane array-FTIR (FPA-FTIR) microscopy, offer a contact-less FTIR solution, reaching LoDs as low as 450 pg when used with a cryogenically cooled MCT array. [26] However, diffraction limits the reliable analysis of single particles with sizes below 10 m. [23, 24]
@@
Quantum cascade laser infrared (QCL-IR) spectroscopy provides higher sensitivity, albeit with a narrower spectral range and additional issues with coherence artifacts [23]. Although QCLs cover the fingerprint region relevant for polymer identification and provide stronger spectral signals due to their high source intensity, they typically have higher noise levels and limited spectral coverage, missing informative regions such as the C–H stretching vibrations around 3000 cm-1 and the low-wavenumber region around 700 cm-1. Full-range spectra, as enabled by FTIR, facilitates chemometric analysis and spectral deconvolution, which can help identify and quantify nanoplastics in multicomponent samples. While such methods can also work within narrower spectral windows, having access to broader spectral information can make differentiation easier and more reliable. QCL-IR microscopy is a complementary method to FTIR microscopy and is therefore sometimes offered in tandem.222HYPERION II from Bruker Corporation Primpke et al. [27] demonstrated that QCL-based microscopy can detect polymer particles as small as 1.4 m. All standard IR spectroscopy techniques suffer from Mie scattering-related spectral artifacts when measuring particulate samples. [28]
@@
Scattering, diffraction, and coherence issues can be reduced by QCL-based photothermal scanning methods, such as optical–photothermal IR (O-PTIR) spectroscopy or AFM-IR microscopy. O-PTIR has proven effective for the imaging with sub-micrometer lateral resolution of individual PS nanoparticles as small as 250 nm in mammalian tissues [29] and identification of plastics (> 600 nm) containing polysiloxanes and imides, originating from the steam disinfection of silicone baby teats. [30] However, O-PTIR is sensitive to the relative focus alignment between the pump and probe beams, and to the axial signal distribution, which can introduce artifacts in depth-resolved measurements.[31] Even better lateral resolution and single nanoparticle analysis can be obtained by AFM-IR. [32, 23] However, AFM-IR faces challenges with low signal-to-noise ratio (SNR) and slow imaging speed. [23]
@@
Raman spectroscopy has the advantage of probing at visible wavelengths with low interference from water and, similar to O-PTIR, allows for the analysis of plastic particles down to sub-micrometer sizes, with spatial resolution sufficient to resolve and chemically identify individual particles. [33] Surface-enhanced Raman spectroscopy (SERS) has emerged as a promising technique. Zhou et al. [34] and Hu et al. [35] reported LoDs as low as 5 g/mL for PS particles as small as 50 nm using conventional SERS substrates. Hyperspectral stimulated Raman scattering (SRS) microscopy has been demonstrated as a powerful tool for the chemical imaging of nanoplastics, enabling identification and classification of particles down to 130 nm. [36] However, SERS suffers from issues such as spectral artifacts and reliance on carefully engineered substrates, limiting its versatility. [11, 37] SRS microscopy, despite offering rapid detection and high sensitivity, relies on complex and expensive instrumentation. [38] Table LABEL:tab:Comparison_between_analytical_techniques summarizes the advantages and disadvantages of the most commonly used analytical techniques for nanoplastic identification.
@@
Although existing analytical techniques such as Py-GC/MS, QCL-IR, O-PTIR, or SERS offer high sensitivity and specificity, they remain impractical for routine nanoplastics monitoring, particularly for water suppliers and official control laboratories. [39]
@@
Nanoelectromechanical systems IR spectroscopy (NEMS-IR) is a technique based on the photothermal effect that has been introduced in recent years. It employs NEMS resonators, or chips, as its key component. [40, 41] The NEMS chip can be used simultaneously as a sample carrier and detector. The NEMS chips, typically consisting of a pre-stressed few-nanometer-thin silicon nitride (SiN) membrane, are robust and allow for a wide variety of sampling techniques, [42] including direct aerosol collection [43, 41, 40] and drop casting. When a sample deposited on the NEMS chip’s surface absorbs IR light, local heating occurs, causing thermal expansion and tensile stress reduction. The corresponding frequency detuning of the NEMS chip is proportional to the absorbed power, and the frequency shift can be monitored using a closed-loop oscillation scheme in combination with a frequency counter, as demonstrated by Bešić et al. [44, 45] Unlike in imaging techniques, the entire sensing area of the NEMS chip is illuminated by the probing IR light. The resulting IR spectrum reveals the bulk properties of the sample. NEMS-IR has shown picogram sensitivity at room temperature over a large spectral range, from the ultraviolet to the far-IR. [42] This technique has been used to analyze polymer nanoparticles, [46, 47] pharmaceutical compounds, [41, 48] polymer micelles, [49] thin polymer films [25] and explosives. [50] NEMS-IR with in-situ thermal desorption (TD) analysis has also been demonstrated for separating simple analyte mixtures by Luhmann et al. [40]
@@
NEMS-IR has only been used for dispersive spectroscopy, typically relying on narrow-band tunable QCLs. Optomechanical metalized cantilevers have previously been used as photothermal detectors in combination with FTIR sources, [51] but their lower sensitivity, [52] reliance on optical readout, and limited surface area for sample collection have constrained their practical application. This paper introduces NEMS-FTIR spectroscopy, interfacing NEMS with FTIR spectrometers as a light source (see Fig. 1A).
By combining the broad spectral coverage, wide availability, and ease of use of commercially available FTIR spectrometers with the high sensitivity of nanomechanical photothermal detection using NEMS resonators, NEMS–FTIR enables measurements with picogram-level detection limits – without the need for cryogenic cooling. Because the readout is purely photothermal, NEMS–FTIR is inherently immune to common IR spectral artifacts, including Mie-scattering distortions [53], ATR-related spectral anomalies [54, 55, 56, 57, 58], and coherence issues.[59]
The responsivity of the NEMS chip’s membrane is highest at its center, where a circular microperforation is located (see Fig. 1B). Consequently, the signal is maximized when the deposited analyte is confined to this region. Previous studies have achieved such localization using an aerosol-based deposition method.[43, 41] In this approach, liquid samples are nebulized and the resulting aerosol particles, containing the analyte, are drawn through the central perforation of the NEMS chip, where they are collected with high efficiency. [40, 41] Although this method yields a uniform coverage of the entire perforated area (see Fig. S1B), the particle-size-dependent collection efficiency complicates quantitative analysis. To address this limitation, we introduce two drop-casting deposition methods that ensure 100% retention of non-volatile analytes, thereby enabling quantitative measurements.
Combined with these quantitative deposition strategies, the straightforward sampling of nanoparticle dispersions directly onto the disposable NEMS chip makes NEMS–FTIR a powerful bulk analysis method for the routine detection and characterization of nanoplastics.
| Technique | Advantages | Disadvantages | References |
| Chemical imaging techniques | |||
| µ-FTIR / FPA-FTIR microscopy |
Broad spectral range
450 pg LoD with cooled MCT array |
Diffraction-limited; unreliable below 10 m particles
High instrumentation cost |
[23, 24, 26] |
| QCL-IR microscopy |
High signal-to-noise ratio
Detects particles down to 1.4 m |
Narrow spectral range
Higher noise than FTIR Coherence artifacts High instrumentation cost |
[23, 27] |
| O-PTIR |
Avoids Mie/diffraction artifacts
High signal-to-noise ratio Sub-micron IR imaging down to 250 nm |
Narrow spectral range
Long acquisition times High instrumentation cost |
[29, 30, 23] |
| AFM-IR |
High spatial resolution (20 nm)
Effective for particles from 20 nm up to 1 m |
Narrow spectral range
Long acquisition times Low signal-to-noise ratio |
[32, 23, 10, 11, 24] |
| Raman microscopy |
High spatial resolution
Low water interference Suitable for particles 1-100 m |
Fluorescence interference
Long acquisition times Low sensitivity for some polymers |
[11, 10, 12, 9] |
| SERS microscopy |
Very high sensitivity
Detects particles down to 50 nm 5 g/mL LoD |
Requires engineered substrates
Long acquisition times Spectral artifacts Polymer-dependent enhancement |
[34, 35, 60, 11, 37] |
| SRS microscopy |
Rapid imaging with high contrast
Detects particles down to 130 nm |
High instrumentation cost | [36, 38, 24] |
| Bulk techniques | |||
| Py-GC/MS |
Ability to detect polymers and plastic aditives
Applicable to complex matrices 1-10 ng LoDs |
Destructive technique
Reduced chemical specificity for micro- and nanoplastics No standardized protocols Background interference Small injection volume High instrumentation cost |
[16, 18, 11, 12, 10, 61, 13, 14, 17, 24, 15, 22] |
| TD-PTR-MS |
High sensitivity for small sample sizes
340 pg LoD |
Destructive technique
Complex data analysis High instrumentation cost |
[11, 20, 12] |
| ATR-FTIR |
Widely available and easy to use
10 ng LoD |
Physical contact with the sample
Risk of cross-contamination Prone to spectral artifacts |
[23, 24, 25, 9] |
| NEMS-FTIR |
No lower particle-size limit
Minimal spectral artifacts Compatible with transmission FTIR spectral libraries Allows downstream analysis by complementary methods such as SEM and O-PTIR Intrinsic SiN internal standard 101-353 pg LoD |
Requires disposable sensing chips
Upper particle-size limit (circa 5 m) Measurements under vacuum preclude the analysis of volatile compounds |
This work |
@@
Many studies have analyzed the release of micro- and nanoplastic particles from teabags during tea preparation using different analytical methods such as scanning transmission X-ray microscopy, [62] IR photothermal heterodyne microscopy, [63] Raman imaging, [64, 65, 66, 67], liquid chromatography-MS-based techniques, [68, 67, 69], and ATR-FTIR spectroscopy. [70, 71] To avoid the complexity of the tea leaves matrix, these investigations were commonly done with emptied teabags, which simplifies sample preparation [72, 63, 64, 70, 71, 68]. Several experimental strategies deviate notably from everyday brewing conditions, for example by using much higher teabag-to-water ratios [71, 70] or involving sample up-concentration steps such as ultracentrifugation, [70] density separation and concentration by ultrafiltration, [62] or evaporation drying to obtain a powder. [71].
In contrast, we demonstrate here that NEMS–FTIR can chemically identify nylon-based PA leachates in the brewing water from a single teabag. NEMS-FTIR analysis of the leachate required just a few nL sample aliquot and did not require any pre-concentration steps, even in the presence of the organic tea leaves matrix. The absence of size limitations in NEMS-FTIR allowed for the identification of both nylon polymer fragments as well as smaller nylon oligomers leached from the nylon teabags. Morphological information could be obtained by imaging samples deposited on NEMS chips directly by scanning electron microscopy (SEM). The LoD of NEMS-FTIR was determined for three model nanoplastics – PS, PP, and PVC – and the measured absorbance was quantitatively related to the actual mass of PS particles deposited on the NEMS chips.
@@
Drop casting of aqueous nanoplastics dispersions on the NEMS chips
To enable quantitative measurements, two drop casting methods were used in this work to deposit and confine nanoplastic samples from a water dispersion at the perforated center of the NEMS chip. 1) Nanoliter droplets were deposited with a piezoelectric nanodroplet dispenser. Following a 2-minute drying time for a 20 nL droplet, the particles were fully confined within the perforated area of the NEMS chips, as shown in Fig. 1C. The coffee ring diameter (see Fig. S2), corresponding to the dried droplet footprint on the membrane, was determined from six independent measurements, yielding an average size of 25223 m for 20 nL droplets, confirming the reproducibility of droplet size. 2) For very low concentration samples, droplets as large as 500 nL were deposited on the NEMS chip. These droplets dispensed via a micropipette have a diameter larger than the 600 m diameter of the perforated area of the membrane. Drying using a combination of permeation and evaporation, or pervaporation, as schematically shown in Fig. 1D, was used to collect the analytes in the sensing area. During the pervaporation process, a controlled humidity gradient was created across the membrane to drive solvent evaporation preferentially through the membrane perforation. As a result, solutes and particulates deposited from large droplets consistently dried in the center of the membrane, as shown in Fig. 1E. By contrast, without pervaporation, the analyte dried across the entire chip (see Fig. 1F). After drying, further droplets can be added to concentrate the sample. A 500 nL drop dried within approximately 30 minutes using the pervaporation method.
@@
Aqueous dispersions of three different polymers, PS (nominal diameter: 100 nm), PP (nominal diameter: 54 nm), and PVC (nominal diameter: 262 nm), were deposited on individual NEMS chips using piezoelectric nanodroplet dispenser. Fig. 2A shows PS nanoparticles after deposition onto the central circular perforation of the NEMS chips (see Fig. 1B). Corresponding scanning electron microscopy (SEM) images of PP and PVC nanoparticles deposited on NEMS chips are provided in the Supporting Information in Fig. S3. These SEM images highlight the generally spherical morphology and relatively uniform size distributions of the model nanoplastic particles studied.
Spectral data were obtained by depositing a 20 nL drop of various model nanoplastic concentrations on the NEMS chips (N=3 for each concentration). Fig. 2B presents a comparison between the NEMS-FTIR spectrum of 20 ng of PS, expressed in absorbance, and a reference spectrum calculated from refractive index data [73] for an equal sample mass. Comparing the absorbance values of the two spectra enabled the calibration of the NEMS-FTIR method, determining , a model parameter representing the effective IR beam diameter illuminating the NEMS membrane. is specific to the IR-light source. The extracted value, mm, is smaller than the theoretical maximum of 1 mm set by the SiN membrane diameter, assuming a beam with a uniform intensity profile. This reduction indicates that the irradiance profile of the IR beam was highest at the center of the membrane, where the sample is located. The details of the spectral processing procedure are described in "Processing of NEMS-FTIR spectra" in the Materials and Methods section.
In the NEMS-FTIR set-up, IR light from the source enters the vacuum chamber from below and passes through the resonator membrane before reaching the sample on its top surface (see inset in Fig. 1A). The resonator membrane exhibits a broad SiN absorption band with a maximum at 835 cm-1 (see Fig. S4). This feature can be exploited as an intrinsic internal standard to monitor chip-to-chip variability and measurement conditions. Because the SiN band absorbs approximately 20% of the incident IR intensity at its maximum, sample vibrational modes that spectrally overlap with this region may, in principle, appear attenuated, as a fraction of the photon flux is already absorbed by the membrane. Such an effect, therefore, constitutes a potential measurement artifact that should be considered when interpreting spectra in this wavenumber range. Conversely, spectral overlaps could impact the internal standardization procedure. In the present measurements, however, no sample features overlap with the SiN absorption band (see Fig. 2B) and no peak attenuation could be observed. As the spectral region around 835 cm-1 is generally sparsely populated, this potential artifact is typically not an issue.
The NEMS-FTIR spectra of PS nanoplastic particles for each deposited mass, power-corrected, normalized using the SiN peak at 835 cm-1, and converted into absorbance, are shown in Fig. 2C. The spectra of PP and PVC nanoparticles are available in the Supporting Information in Figs. S5 and S6, respectively.
@@
The IR peak positions in the recorded NEMS-FTIR spectra of PS nanoparticles are in agreement with data reported in the literature. [74, 75, 76, 77, 78] Notably, PS is characterized by peaks at 753 and 700 , as well as by a mono-substituted benzene ring, as evidenced by a series of weak overtone and combination bands between 2000 and 1650 (so-called ’benzene fingers’), often difficult to detect in the spectrum if the concentration of molecules containing the benzene ring is low. [76] However, these peaks are clearly visible in the NEMS-FTIR spectra of PS, even for the very low masses in the nanogram range. The exact peak positions are listed in Table S1 in the Supporting Information.
@@
The mass of PS nanoparticles present on the NEMS chips was estimated using the attenuation coefficient of PS and the absorptances calculated from the NEMS-FTIR spectra at the wavenumber of . The mass estimation procedure is detailed in the "Mass estimation" in the Materials and Methods section. The calculation of the sample absorptance accounted for the spatial distribution of the particles, which predominantly formed a circular patch in the center of the perforation on the NEMS chips. Fig. S7 shows the correlation between the deposited (amount of deposited PS particles) and estimated (from the measured absorptances of PS particles) mass of nanoplastic particles. The uncertainty in the sample mass deposited on the NEMS chip was estimated from the uncertainty in the tools used for sample preparation and for sample deposition. The largest uncertainty on the sample mass as estimated from the NEMS-FTIR spectra arises from the variability in the measured absorptance . Exact values for the masses deposited on the NEMS chips and estimated masses as calculated from the NEMS-FTIR spectra, along with their standard deviations, can be found in Table S2 in the Supporting Information.